OpenStudy (loser66):

Let \(f:\mathbb R \rightarrow \mathbb R\) be \(C^1\) and \(u=f(x) \), \(v= -y+xf(x)\) If \(f'(x_0) \neq 0\), show that this transformation is invertible near \(x_0,y_0)\) and the inverse has the form \(x= f^{-1}(u)\) and \(y= -v+uf^{-1}(u)\) Please, help

OpenStudy (lilshane):

what you think it is

OpenStudy (jordan123321):

hi @lilshane

OpenStudy (loser66):

It is a "show" problem, It doesn't have a result

OpenStudy (loser66):

My attempt: The matrix of differential is \[Df(u,v)_{(x,y)} =\left[\begin{matrix}\dfrac{\partial u}{\partial x}&\dfrac{\partial u}{\partial y}\\dfrac{\partial u}{\partial y}&\dfrac{\partial v}{\partial y}\end{matrix}\right]=\left[\begin{matrix}f'(x)&0\\f(x)+xf'(x) &-1\end{matrix}\right]\]

OpenStudy (loser66):

det (Df) = \(-f'(x) \) , and given condition gives us \(f'(x_0)\neq 0\), hence the function is invertible.

OpenStudy (loser66):

and the inverse matrix of D is \[\left[\begin{matrix}-1&0\\-f(x) -xf'(x) &f'(x)\end{matrix}\right]\]

OpenStudy (loser66):

My argument goes wrong from now :(

OpenStudy (loser66):

That is \(\dfrac{\partial x}{\partial u}=-1\), hence x = -u +C but from the original one, u = f(x) hence \(x= f^{-1}(u)\) that is the correct one. However, I don't know what is wrong with my argument.

OpenStudy (loser66):

@bibby

OpenStudy (loser66):

@Kainui

OpenStudy (loser66):

@rational

OpenStudy (anonymous):

I guess you are answering perfectly your quest :)